OpenStudy (anonymous):

Does 2/3 +3/2 = ((2+3))/((3+2) )=1 ? Explain your answer

OpenStudy (fateal):

try algebra calculator?

OpenStudy (fateal):

nvm...ima look it up n tell u x3 one sec plz

OpenStudy (fateal):

2/3 +3/2 =2.166667 ((2+3))/((3+2) )= 1 so the answer is no? theyre not eqaul

OpenStudy (fateal):

did that help? i just went to http://www.mathpapa.com/algebra-calculator.html then i plugged in the first number then the second

OpenStudy (fateal):

glad :) can i get a medal? lel o.e

OpenStudy (fateal):

ty ^^